OpenStudy (anonymous):

solve each equation: 5x^2 - 2x + 1 = 0 7x^2 + 8x = -6

OpenStudy (abb0t):

Use quadratic equation for the first one: \[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] For the second one, add the 6 onto the equation to make it in the form: \[ax^2 + bx + c = 0\]

OpenStudy (anonymous):

Thanks abb0t, I figured it out tho

OpenStudy (anonymous):

The first has roots that are complex ie. .2+0.4i and .2-.4i